Pin at Manda Method of Joint Find all members oma 16


Problem 2 6 @ 12 ft = 72 ft. 16 ft 15 k 50 k 40 k 30k 20° B :201- 20 k 10 k
Answer #1

20 20 20.8 Sine = 16 16 +12 Caso - 12 = tano - 0.6 Joint : Fay As no force acting at G. Fara Far = 0 Joint (It) (of 20° L 4Joint E. fax Sino = 0.8k Iffe= 1k ffic Case +FFE = 0 IFFE 2-0.6k Joint Co- I Fius - FKL tffie Calzo Fiks = 0.6k + (2] (0.6k FJoint I fjk FJI = fjke + FJE Caso FJI = 1.2k + 2 (0.6k IFJF = 1.8k Fre find top = 0 Fop-dok Fos Joint D - FOI A oto Fe for fiJoint © 8 oto Т А o teft case the fee Fun (0.6) then = -1.8k I for a link Felp fino thos- I fan = -1k ATAU CD hi part Joint (Force in members is determined by equilibrium equations at every joints.